OpenStudy (anonymous):

Find the least common multiple of x³ – x² + x – 1 and x² – 1 . Write the answer in factored form.

OpenStudy (luigi0210):

Factor both and what do you get?

OpenStudy (anonymous):

(x - 1)(x - 1)(x + 1) and (x - 1)(x + 1)?

OpenStudy (anonymous):

(x - 1)(x - 1)(x + 1) and (x - 1)(x + 1)?

OpenStudy (anonymous):

Can I just like say the answer?

OpenStudy (anonymous):

Do you know how to solve it shelby?

OpenStudy (luigi0210):

Close.. it would actually be: \[\LARGE (x^2+1)(x-1)~and~(x+1)(x-1)\]

OpenStudy (anonymous):

(x + 1)(x – 1)(x² + 1)

OpenStudy (anonymous):

How did you get that @Anonymouseee

OpenStudy (anonymous):

x^3 - x^2 + x - 1 = (x - 1)(x^2 + 1) That is the big key to the problem and that was done using synthetic division. x^2 - 1 = (x - 1)(x + 1) So, you take (x - 1)(x^2 + 1) from the first, but you only have to take "x + 1" from the second because you already have a "x - 1" factor from the first. And that's it!

OpenStudy (luigi0210):

Yea, factor an x^2 and 1.. \[\LARGE x^2(x-1)+1(x-1)\] \[\LARGE (x^2+1)(x-1)\]

OpenStudy (anonymous):

nuff siad.

OpenStudy (anonymous):

said**

OpenStudy (anonymous):

So...... Did it help?

OpenStudy (anonymous):

I see thanks

OpenStudy (anonymous):

No problem